OpenStudy (anonymous):

Write the equation of a line that passes through the points (8, 5) and (11, 14).

OpenStudy (anonymous):

this is what i got y – 5 = 3(x – 8)

OpenStudy (anonymous):

ok solve it now

OpenStudy (anonymous):

that is solved

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

y + 5 = 3(x + 8) y – 5 = 3(x – 8) y + 5 = (x + 8) y – 5 = (x – 8)

OpenStudy (anonymous):

those are the only answers

OpenStudy (anonymous):

if you have y-5 =3(x+8), put it in y=mx+b form

OpenStudy (anonymous):

y-5 = 3x+24 => y=3x+29

OpenStudy (anonymous):

okay but that wasent any of my choices

OpenStudy (anonymous):

Ah I see what you mean.

OpenStudy (wolf1728):

iambatman - recheck your addition

OpenStudy (anonymous):

so was i right?

OpenStudy (anonymous):

Nice catch wolf :)

OpenStudy (wolf1728):

thanks :-)

OpenStudy (anonymous):

Looked at the wrong equation lol.

OpenStudy (wolf1728):

Iambatman posted y-5 = 3x+24 => y=3x+29 but as you can see the equation becomes y = 3x + 19

OpenStudy (anonymous):

y + 5 = 3(x + 8) y – 5 = 3(x – 8) y + 5 = (x + 8) y – 5 = (x – 8)

OpenStudy (anonymous):

so which one?

OpenStudy (anonymous):

Wolf I looked at this first equation y+5=2(x+8) but yes you're correct reliver it's B

OpenStudy (anonymous):

3*

OpenStudy (wolf1728):

Okay, so it is solved right?